# LayeredMAPF: a decomposition of MAPF instance without compromising solvability

### v1

@@ III-B Theory basis of decomposition
To reduce the time cost of searching for paths, we propose a connectivity graph $\mathcal{G}_{c}$, whose nodes consist of free grid groups and the starts and targets of agents, and edges represent connectivity between nodes.
Definition 1
A free grid group is defined as a set of passable grids that contains no start or target grid of agents, and there always exists a path connecting any two grids within the group (where each waypoint in the path is within the same free grid group). We denote a free grid group as $f_{g}$ and a set of free grid groups as $F_{g}$. An example about free grid group are shown in Fig. 2.
A: a MAPF intance
B: free grid groups
C: connectivity graph
Figure 2: These figures depict a grid map and an MAPF instance containing three agents (A), free grid groups (B), and the corresponding connectivity graph. In these figures, ”0s, 1s, 2s” denote the starting positions of the three agents, while ”0t, 1t, 2t” represent their targets. Grey grids without labels indicate unpassable grids, while ”0f, 1f” label the grids belonging to two free grid groups, respectively.
Definition 2
The connectivity graph $\mathcal{G}_{c}$ is an undirected graph whose nodes consist of free grid groups, starting points, and targets for all agents, with edges representing the connectivity between these nodes.
Definition 3
@@
As an example, the dependence path for agent 0 in the instance shown in Fig. 2 is ”0s $\rightarrow$ 0f $\rightarrow$ 1t $\rightarrow$ 1f $\rightarrow$ 0t”; for agent 1, it is ”1s $\rightarrow$ 1f $\rightarrow$ 1t”; and for agent 2, it is ”2s $\rightarrow$ 1f $\rightarrow$ 2t”.
@@
1, if the dependence path of agent $a_{i}$ contains only the start of agent $a_{j}$, it indicates that $a_{i}$ should be solved after $a_{j}$;
2, if the dependence path of agent $a_{i}$ contains only the target of agent $a_{j}$, it implies that $a_{i}$ should be solved before $a_{j}$;
@@
It is worth noting that an agent’s dependence path is not unique. By updating their dependence paths, we can alter the relationship between two agents (in implementation, we check for the existence of such paths rather than explicitly maintaining all agents’ dependence paths). By changing the relationship between two agents, we can update the subproblems of a Multi-Agent Path Finding (MAPF) instance’s decomposition. Although there is no direct method to ensure that agents’ dependence paths correspond to smaller subproblems, we have proposed several approaches to encourage the generation of smaller subproblems; their details can be found in Section IV.
Here, we define the function for searching dependence paths as $p_{i}$ = $search\_path$($a_{i}$, $\mathcal{G}_{c}$, $available\_nodes$, $avoid\_nodes$). Here, $\mathcal{G}_{c}$ represents the connectivity graph of the MAPF instance, available nodes are those nodes in $\mathcal{G}_{c}$ that can be traversed, and avoid nodes are those nodes that must be avoided.
Specifically, an empty set for available nodes means that all nodes are available. If the $search\_path$ function succeeds, it returns a path; otherwise, it returns an empty set. If not specified, both available nodes and avoid nodes default to being empty.
In detail, the $search\_path$ function is implemented using a classic Best First Search approach, similar to A*. For this function, the cost and heuristic value represent the number of nodes the path has visited and the number of agents the path will visit. We consider a path to have visited an agent if it contains the start or target of that agent.
IV Methodology
@@
As mentioned previously, we have not identified a direct method to decompose a MAPF instance into multiple subproblems. Thus, we initiate the process with an initial decomposition, progressively refining it to generate smaller subproblems while ensuring that each step maintains solvability. We define two types of subproblems encountered in various stages: clusters, which have no restrictions on the order of solving, and levels, which impose limitations on the order of solving.
@@
1, identifying all agent dependencies and establishing initial clusters based on them;
@@
3, further refining agent dependencies to decompose clusters into levels and sorting;
Step 1 and 2 are described in Section IV-A, and step 3 is described in Section IV-C.
An overview of the decomposition process of a MAPF instance is depicted in Figure 1.
Following decomposition, we explore methods to solve subproblems independently and then combine their results to obtain a conflict-free solution for the original MAPF instance, as detailed in Section IV-D.
IV-A Decomposetion to clusters
@@
For the three dependence paths mentioned before, agent 0 and agent 1 are relevant, while agent 2 is irrelevant to either of them.
Definition 5
@@
Definition 6
Cluster: a maximal connected subgraph (i.e., connected component) of $\mathcal{G}_{a}$.
Intuitively, a cluster is a set of agents from the raw MAPF instance, where any agents within it are only relevant to agents within the same set (based on their dependence paths). A cluster is denoted as $r$, and a set of clusters is denoted as $\mathcal{R}$. Additionally, we refer to an agent set that constitutes a cluster as an independent agent set.
For the three dependence paths mentioned before, agent 0 and agent 1 form one cluster, while agent 2 forms another cluster.
An intuitive way to obtain clusters is by extracting clusters from the dependence paths that visit the fewest number of agents for each agent. We refer to such clusters as initial clusters.
@@
There is room for decomposing these initial clusters into smaller ones by updating the dependence paths of agents (in implementation, we verify the existence of such paths rather than explicitly maintaining all agents’ dependence paths). Thus, we propose a method to iteratively bipartition initial clusters until further subdivision is not possible, aiming to minimize the size of subproblems. More details about bipartitioning can be found in the following section. This process ensures the solvability of decomposition, and the more bipartitioning steps, the better the results obtained, although it does not guarantee the discovery of the optimal decomposition. An overview of the process of decomposing an instance into clusters is outlined in Algorithm 1.
Input: $\mathcal{C}_{\mathcal{N}},\mathcal{G}_{c}$
Output: $\mathcal{R}$
1 ///// / determine initial clusters
2 $P=\{p_{1},p_{2},...,p_{k}\}$; ///// / dependence path of all agents
3 for i = 1, 2,…,$k$ do
4       $p_{i}=search\_path(a_{i},\mathcal{G}_{c},\emptyset,\emptyset,\emptyset)$;
5
6get $\mathcal{G}_{a}$ from $P$;
7 ///// / get initial clusters
8 $\mathcal{R}_{init}$ = all connected components of $\mathcal{G}_{a}$;
9 $\mathcal{R}=\emptyset$; ///// / final clusters
10 for $r\in\mathcal{R}_{init}$  do
11       $temp\_cluster=r$;
12       while $temp\_cluster\neq\emptyset$ do
13             $\{r_{major},r_{remain}\}$ = Bipartition of cluster ($r,\mathcal{G}_{c}$);
14             add $r_{major}$ to $\mathcal{R}$;
15             $temp\_cluster=r_{remain}$;
16
17
return $\mathcal{R}$;
@@ 15             $temp\_cluster=r_{remain}$;
IV-B Bipartition of clusters
@@
Input: $r,\mathcal{G}_{c}$
Output: $\{r_{major},r_{remain}\}$
1 ///// / step 1
2 $\mathcal{G}_{u}=\emptyset$;
3 for $a_{i}\in r$ do
4       for $a_{j}\in\$r$ do
5             if $a_{i}\neq a_{j}$ then
6                   if $search\_path(a_{i},\mathcal{G}_{c},r,{a_{j}},\emptyset)=\emptyset$  then
7                         add $(a_{i}\rightarrow a_{j})$ as an edge to $\mathcal{G}_{u}$;
8
9
10
11
[…]
13 $r_{remain}=r-r_{major}$;
14 while $True$ do
15       ///// / step 2
16       $new\_agents=\emptyset$;
17       while $True$ do
18             for $a_{i}\in r_{remain}$ do
19                   if $search\_path(a_{i},\mathcal{G}_{c},r,r_{remain},\emptyset)=\emptyset$ then
20                         move $a_{i}$ from $r_{remain}$ to $r_{major}$;
21                         add $a_{i}$ to $new\_agents$;
22
23
24            if $r_{remain}$ is independent $or$ $r_{remain}$ is empty then
25                   break;
26
27      ///// / step 3
28       while $True$ do
29             for $a_{j}\in new\_agents$ do
30                   if $search\_path(a_{i},\mathcal{G}_{c},r_{major},r_{remain})=\emptyset$ then
31                         ///// / consider $a_{i}$ belong to cluster $r$, following search path always success $p_{i}=search\_path(a_{i},\mathcal{G}_{c},r,\emptyset,r\_major)$;
32                         move agents in ($p_{i}\cap r_{remain}$) from $r_{remain}$ to $r_{major}$;
33
34
35            if $r_{major}$ is independent then
36                   break;
37
38      ///// / step 4
39       if both $r_{major}$ and $r_{remain}$ are independent then
40             break;
41
return $\{r_{major},r_{remain}\}$;
@@ 16       $new\_agents=\emptyset$;
Definition 7
A crucial concept to introduce is the notion of ”unavoidable agents” for an agent $a$ within a cluster $r$. Unavoidable agents of an agent $a$ in a cluster $r$ refer to those agents within $r$ that must be traversed by $a$’s dependence path. In other words, these agents must belong to the same cluster and cannot be further divided.
Definition 8
Unavoidance graph $\mathcal{G}_{u}$ of a cluster: an undirected representation of whether one agent is unavoidable to another agent within the cluster. It is important to note that for a given cluster $r$, its unavoidance graph is unique.
@@
Intuitively, these maximum unavoidable agents represent the largest undividable subset within the current cluster, are referred to as ”major set” during the cluster bipartition process. Agents within the cluster, excluding those belonging to the maximum unavoidable agents, are referred to as ”remaining set” during the cluster bipartition process.
It is noteworthy that during the bipartition process, agents in the remaining set will be moved to the major set.
Input: $r,\mathcal{G}_{c}$
Output: $\mathcal{L}$
1 ///// / determine initial dependence paths
2 $P=\emptyset$; ///// / $P[a]$ means agent a’s denpendence path
3 for agent $a$ in $r$ do
4       $p_{i}=search\_path(a_{i},\mathcal{G}_{c},r,{a_{j}})$;
5       add $p_{i}$ to $P$;
6
7get $\mathcal{G}_{s}$ from $P$;
8 determine all loop in $\mathcal{G}_{s}$ via strong component detect;
9 ///// / sort levels;
10 construct graph $\mathcal{G}_{l}$ via $\mathcal{G}_{s}$;
11 $root\_levels$ = level in $\mathcal{G}_{l}$ that not later than other level;
12 set order of all level to 0;
13 $levels$ = $root\_levels$;
14 while levels is not empty do
15       $next\_levels$ = $\emptyset$;
16       for level in levels do
17             for another_level that later than level in $\mathcal{G}_{l}$ do
18                   if $another\_level$’s order $\leq$ level’s order  then
19                         $another\_level$’s order = level’s order + 1;
20                         add $another\_level$ to $next\_levels$;
21
22
23
24      $levels$ = $next\_levels$;
25return all levels in the order of last visit to them;
@@ 24      $levels$ = $next\_levels$;
1, identify the maximum unavoidable agents of the cluster, referred to as the major set;
2, examine each agent in the remaining set to determine if its dependence path must pass through agents of the major set. Move agents meeting this criterion to the major set;
3, Check any newly added agents to the major set to ascertain if their dependence paths must pass through agents in the remaining set. Transfer such agents from agents the remaining set to the major set;
4, verify whether the major set and the remaining set both meet the requirements of the cluster. If they do (or if the remaining set is empty), exit and return the major set and the remaining set as the result of the bipartition. Otherwise, proceed to step 2;
@@
The pseudocode for bipartitioning a cluster is outlined in Algorithm 2.
Input: $\{P_{1},...,P_{m}\}$ ///// / assume there are $m$ subproblems and their solution is $P_{i}$
Output: $\{P_{1},...,P_{m}\}$
1 set last occupied time of all grids to 0;
2 for each solution $P$ in $\{P_{1},...,P_{m}\}$ do
3       while True do
4             $t$=1; ///// / current time index
5             ///// / check whether insert wait action to $P$
6             $need\_wait$ = False;
7             $all\_finished$ = True;
8             for path $p$ in $P$ do
9                   if t $>$ size of $p-1$ then
10                         ///// / don’t check path that have reach target
11                         continue;
12
13                  else
14                         $all\_finished$ = False;
15
16                  ///// / $p[t]$ means path $p$’s location at time $t$
17                   if last occupied time of $p[t]\geq t$ then
18                         $need\_wait$ = True;
19                         break;
20
21
22            if all_finished then
23                   break;
24            ///// / insert wait action to all path in $P$
25             if need_wait then
26                   for path $p$ in $P$ do
27                         if t $>$ size of $p-1$ then
28                               continue;
29
30                        repeat $p[t]$ in $p$ at time index $t$;
31
32
33            else
34                   ///// / update last occupied time of grids
35                   for path $p$ in $P$ do
36                         if t $>$ size of $p-1$ then
37                               continue;
38
39                        set last occupied time of $p[t]$ to $t$;
40
41
42            $t$ = $t$ + 1;
43
44
45return $\{P_{1},...,P_{m}\}$;
@@ 42            $t$ = $t$ + 1;
Input: $\mathcal{C}_{\mathcal{N}},A$, MAPF
Output: $P$
1 ///// / 1, decomposition of MAPF instance
2 construct $\mathcal{G}_{c}$ from $\mathcal{C}_{\mathcal{N}},A$;
3 $R$ = Decomposing to cluster($\mathcal{G}_{c}$);
4 $L=\emptyset$;
5 for cluster $r$ in $R$ do
6       $L^{\prime}$ = Decomposing to leve and sorting($r$, $\mathcal{G}_{c}$);
7       merge $L^{\prime}$ to $L$;
8
[…]
10 ///// / 2, solve subproblems
11 for $l$ in $L$ do
12       if MAPF is serial then
13             set later subproblem’s start in $\mathcal{C}_{\mathcal{N}}$ to unpassable;
14             $current\_solution$ = MAPF($\mathcal{C}_{\mathcal{N}},l,all\_subproblem\_paths$);
15             add $current\_solution$ to $all\_subproblem\_paths$;
16             reset $\mathcal{C}_{\mathcal{N}}$;
17
18      else
19             set previous subproblem’s target and later subproblem’s start in $\mathcal{C}_{\mathcal{N}}$ to unpassable;
20             $current\_solution$ = MAPF($\mathcal{C}_{\mathcal{N}},l$);
21             add $current\_solution$ to $all\_subproblem\_paths$;
22             reset $\mathcal{C}_{\mathcal{N}}$;
23
24
25///// / 3, merge solutions of subproblems
26 if MAPF is serial then
27       $P$ = $all\_subproblem\_paths$;
28
29else
30       $P$ = Merge results($all\_subproblem\_paths$);
31
32return $P$;

### v2

@@ 3.2.2 Connectivity graph
We then propose a connectivity graph, $\mathcal{G}_{c}$, whose nodes consist of free cell groups and the start and target states of agents, with edges representing the connectivity between these nodes. Since $\mathcal{G}c$ has a smaller size than $\mathcal{C}\mathcal{N}$, the time cost of searching for paths can be significantly reduced, especially for instances with sparse agents.
In the implementation, since free cell groups do not influence whether an agent’s dependent path passes through other agents’ states, we simplify $\mathcal{G}_{c}$ by directly connecting the nodes representing agent states if they can be connected via free cell groups. After simplification, $\mathcal{G}c$ has $2k$ nodes, where $k$ is the number of agents. Compared to $\mathcal{C}\mathcal{N}$, the simplified $\mathcal{G}_{c}$ has a significantly smaller size, which accelerates the process of determining whether two agents can be in different subproblems.
An example of the connectivity graph is shown in Fig. 8.
A
B
C
Figure 8: These figures show the free cell groups (marked as $F_{1},F_{2},\dots$, in Figure A), the connectivity graph $\mathcal{G}_{c}$ (in Figure B), and the simplified connectivity graph (in Figure C) of the instance in Fig. 2 A.
3.2.3 Variants of search path
To reduce the time cost of searching for paths, we search for paths in $\mathcal{G}c$ rather than in $\mathcal{C}\mathcal{N}$. We denote the paths in $\mathcal{G}_{c}$ as dependence paths.
Definition 4.
@@ 3.2.3 Variants of search path
As an example, the dependence path for agent 1 in the instance shown in Fig. 8 is “$S_{1}\rightarrow T_{2}\rightarrow F_{1}\rightarrow T_{3}\rightarrow T_{1}$"; for agent 2, its dependence path is “$S_{2}\rightarrow F_{2}\rightarrow T_{3}\rightarrow F_{1}\rightarrow T_{2}$"; and for agent 3, the dependence path is “$S_{3}\rightarrow F_{1}\rightarrow T_{3}$".
@@
1, if the dependence path of agent $a_{i}$ contains only the start of agent $a_{j}$, it indicates that $a_{i}$ should be solved after $a_{j}$ to enable they could be in different subproblems;
2, if the dependence path of agent $a_{i}$ contains only the target of agent $a_{j}$, it implies that $a_{i}$ should be solved before $a_{j}$ to enable they could be in different subproblems;
@@
It is worth noting that an agent’s dependence path is not unique. By updating their dependence paths, we can alter the relationship between two agents (in implementation, we check for the existence of such paths rather than explicitly maintaining all agents’ dependence paths). By changing the relationship between two agents, we can update the subproblems of a MAPF instance’s decomposition. Considering that making an agent’s dependence path avoid another agent might cause the path to pass through a third agent, and thus fail to generate smaller subproblems, we have no direct method to ensure that agents’ dependence paths correspond to smaller subproblems, we have proposed several approaches to encourage the generation of smaller subproblems; their details can be found in Section 4.
Here, we define two variants of the search path function used for decomposing the LA-MAPF instance into clusters:
Definition 5.
$search\_path\_agent(a_{i},\mathcal{G}_{c},available\_agents,avoid\_agents)$, where $available\_agents$ represents agents whose start or target states the dependence path of $a_{i}$ can pass through, and $avoid\_agents$ represents agents whose start or target states the dependence path of $a_{i}$ cannot pass through. If $available\_agents=\emptyset$, it means the path can pass through all nodes in $\mathcal{G}_{c}$, except those that belong to the start or target states of agents in $avoid\_agents$.
$search\_path\_agent$ performs a complete graph search on $\mathcal{G}_{c}$ and returns the agents whose dependence paths involve $a_{i}$. If no such dependence path exists, $search\_path\_agent$ returns $\emptyset$.
Similarly, we define a variant of $search\_path$ used in decomposing the instance into levels:
Definition 6.
$search\_path\_SAT(a_{i},\mathcal{G}_{c},available\_SATs,avoid\_SATs)$, where $available\_SATs$ represents the start or target states that the dependence path of $a_{i}$ can pass through, and $avoid\_SATs$ represents the start or target states that the dependence path of $a_{i}$ cannot pass through. If $available\_SATs=\emptyset$, it means the path can pass through all nodes in $\mathcal{G}_{c}$, except those that belong to the start or target states of agents in $avoid_{S}ATs$.
$search\_path\_SAT$ also performs a complete graph search on $\mathcal{G}_{c}$ and returns the start or target states of agents whose dependence paths are involved with $a_{i}$. If no such dependence path exists, $search\_path\_SAT$ returns $\emptyset$.
Here are some examples of the variants of the search path for the instance in Fig. 8:
$search\_path\_agent(a_{1},\mathcal{G}_{c},\{a_{1},a_{2},a_{3}\},\emptyset)=\{a_{1},a_{2},a_{3}\}$;
$search\_path\_agent(a_{1},\mathcal{G}_{c},\emptyset,\{a_{2}\})=\emptyset$ (since $a_{1}$’s dependence path must pass through $T[a_{2}]$);
$search\_path\_agent(a_{1},\mathcal{G}_{c},\{a_{1},a_{3}\},\emptyset)=\emptyset$ (since $a_{1}$’s dependence path must pass through $T[a_{2}]$);
$search\_path\_SAT(a_{1},\mathcal{G}_{c},\emptyset,\emptyset)=\{S[a_{1}],T[a_{1}],T[a_{2}],T[a_{3}]\}$;
$search\_path\_SAT(a_{2},\mathcal{G}_{c},\emptyset,\emptyset)=\{S[a_{2}],T[a_{2}],T[a_{3}]\}$;
$search\_path\_SAT(a_{2},\mathcal{G}_{c},\emptyset,\{T[a_{3}]\})=\emptyset$ (since $a_{2}$’s dependence path must pass through $T[a_{3}]$);
$search\_path\_SAT(a_{2},\mathcal{G}_{c},\{S[a_{2}],T[a_{2}],T[a_{3}]\},\emptyset)=\{S[a_{2}],T[a_{2}],T[a_{3}]\}$;
There are two main advantages of $search\_path\_agent$ and $search\_path\_SAT$ compared to $search\_path$:
  1. 1.
    $\mathcal{G}_{c}$ has a smaller size compared to $\mathcal{C}_{\mathcal{N}}$ (e.g., $\mathcal{C}_{\mathcal{N}}$ in Fig. 8 has 4*4=16 nodes, while the related simplified $\mathcal{G}_{c}$ has only 2*3=6 nodes), so $search\_path\_agent$ and $search\_path\_SAT$ are faster compared to $search\_path$.
  2. 2.
    $search\_path\_agent$ and $search\_path\_SAT$’s return values provide direct information about the current agent’s relationship with other agents, which is very useful in determining whether two agents must be in the same subproblem.
4 Methodology
@@
As mentioned previously, we have not identified a direct method to decompose a MAPF instance into multiple subproblems. Thus, we initiate the process with an initial decomposition, progressively refining it to generate smaller subproblems while ensuring that each step pass the mentioned solvability check. We define two types of subproblems encountered in various stages: clusters, which have no restrictions on the order of solving, and levels, which impose limitations on the order of solving.
@@
1, identifying all agents’ dependence path and establishing initial clusters based on them;
@@
3, further refining agent dependencies to decompose clusters into levels and sorting.
Step 1 and 2 are described in Section 4.1, and step 3 is described in Section 4.3.
An overview of the decomposition process of a MAPF instance is depicted in Fig. 1.
Following decomposition, we explore methods to solve subproblems independently and then combine their results to obtain a conflict-free solution for the original MAPF instance, as detailed in Section 4.4.
4.1 Decomposetion to clusters
@@
Definition 8.
@@
Example about $\mathcal{G}_{a}$ can be found in Fig. 10.
It is important to note that the relevance of two agents is determined by $avail\_agents$, $avoid\_agents$, and $\mathcal{G}_{c}$. While the subgraph $\mathcal{G}_{c}$ is constant, $avail\_agents$ and $avoid\_agents$ are variables, meaning that the relevance between two agents may change as these sets are modified.
Essentially, based on the definition of a cluster, a cluster is a maximal connected component in $\mathcal{G}_{a}$. Intuitively, a cluster is a set of agents from the raw MAPF instance, where any agents within it are only relevant to agents within the same set (based on their dependence paths). Additionally, we refer to an agent set that constitutes a cluster as an independent agent set.
An intuitive way to obtain clusters is by generating clusters from the graph of agent’s relavance determined by the dependence paths that visit the fewest number of agents for each agent. We refer to such clusters as initial clusters(Line 1 $\sim$ 8 in Algorithm 1). An example about initialize of clusters are shown in Fig. 9.
@@
There is room for decomposing these initial clusters into smaller ones by updating the dependence paths of agents. Thus, we propose a method to iteratively bipartition initial clusters until further subdivision is not possible, aiming to minimize the size of subproblems. More details about bipartitioning can be found in the following section. This process ensures the legality of decomposition, and the more bipartitioning steps, the better the results obtained, although it does not guarantee the discovery of the optimal decomposition. An overview of the process of decomposing an instance into clusters is outlined in Algorithm 1.
Figure 9: These figures show the initialization of clusters in a simple instance. Figure A illustrates the instance, where white cells represent passable cells and grey cells represent unpassable cells. The notations “$S_{i}$” and “$T_{i}$” denote the start and target states of agent $a_{i}$.
Figure B displays the simplified connectivity graph of agents.
Figure C shows the relevant agents for each agent, and Figure D presents the graph of agent relevance ($\mathcal{G}_{a}$).
Figure E shows the initial clusters determined by the graph of agent relevance, i.e., all connected components in $\mathcal{G}_{a}$.
Figure 10: These figures demonstrate the process of bipartitioning cluster $C_{1}$, as mentioned in Fig. 9.
Figure A illustrates the unavoidable agents for each agent. Figure B shows the related unavoidance graph $\mathcal{G}_{u}$. Figures C, D, and E depict steps 2, 3, and 4 in the bipartition process of the cluster.
Figure 11: These figures show the verification of three clusters from the instance in Fig. 10. Figure A illustrates the process of decomposing the instance into three clusters, similar to Fig. 1. Figures B, C, and D show the verification of the legality of the three clusters, i.e., checking whether there is a solution for each subproblem’s agents while assuming the other clusters’ agents occupy both their start and target states. Since we ensure that each cluster’s agents are only related to the agents in the current cluster, all clusters are legal. The solutions for each agent are shown as dotted arrowed lines in different colors.
@@
1:$\mathcal{C}_{\mathcal{N}},\mathcal{G}_{c}$
$\mathcal{R}$
3:///// / determine initial clusters
4:$P=\{p_{1},p_{2},...,p_{k}\}$; ///// / dependence path of all agents
5:for i = 1, 2,…,$k$ do
6:     $p_{i}=search\_path\_agent(a_{i},\mathcal{G}_{c},\emptyset,\emptyset)$;
7:end for
8:get $\mathcal{G}_{a}$ from $P$;
9:///// / get initial clusters
10:$\mathcal{R}_{init}$ = all connected components of $\mathcal{G}_{a}$;
$\mathcal{R}=\emptyset$; ///// / final clusters
12:for $r\in\mathcal{R}_{init}$  do
13:     $temp\_cluster=r$;
14:     while $temp\_cluster\neq\emptyset$ do
15:         $\{r_{major},r_{remain}\}$ = Bipartition of cluster ($temp\_cluster,\mathcal{G}_{c}$);
16:         add $r_{major}$ to $\mathcal{R}$;
17:         $temp\_cluster=r_{remain}$;
18:     end while
19:end for
20:return $\mathcal{R}$;
4.2 Bipartition of clusters
@@
1:$r,\mathcal{G}_{c}$
2:$\{r_{major},r_{remain}\}$
3:///// / step 1
4:$\mathcal{G}_{u}=\emptyset$;
5:for $a_{i}\in r$ do
6:     for $a_{j}\in r$ do
7:         if $a_{i}\neq a_{j}$ then
8:              if $search\_path\_agent(a_{i},\mathcal{G}_{c},r,\{a_{j}\})=\emptyset$  then
9:                  add $(a_{i}\rightarrow a_{j})$ as an edge to $\mathcal{G}_{u}$;
10:              end if
11:         end if
12:     end for
13:end for
14:$r_{major}$ = largest connected component of $\mathcal{G}_{u}$;
15:$r_{remain}=r-r_{major}$;
16:while $True$ do
17:     ///// / step 2
18:     $new\_agents=\emptyset$;
19:     while $True$ do
20:         for $a_{i}\in r_{remain}$ do
21:              if $search\_path\_agent(a_{i},\mathcal{G}_{c},r_{remain}$, $r_{major}$ )=)$\emptyset$) = ∅ then
22:                  move $a_{i}$ from $r_{remain}$ to $r_{major}$;
23:                  add $a_{i}$ to $new\_agents$;
24:              end if
25:         end for
26:         if $r_{remain}$ is independent $or$ $r_{remain}$ is empty then
27:              break;
28:         end if
29:     end while
30:     ///// / step 3
31:     while $True$ do
32:         for $a_{j}\in new\_agents$ do
33:              if $search\_path\_agent(a_{i},\mathcal{G}_{c},r_{major},r_{remain})=\emptyset$ then
34:                  ///// / consider $a_{i}$ belong to cluster $r$, following search path always success
35:                  $p_{i}=search\_path\_agent(a_{i},\mathcal{G}_{c},r,\emptyset$);
36:                  move agents in ($p_{i}\cap r_{remain}$) from $r_{remain}$ to $r_{major}$;
37:              end if
38:         end for
39:         if $r_{major}$ is independent then
40:              break;
41:         end if
42:     end while
43:     ///// / step 4
44:     if both $r_{major}$ and $r_{remain}$ are independent then
45:         break;
46:     end if
47:end while
48:return $\{r_{major},r_{remain}\}$;
Definition 9.
Unavoidable agents of an agent $a$ in a cluster $r$ refer to those agents within $r$ that must be traversed by $a$’s dependence path. In other words, these agents must belong to the same cluster and cannot be further divided.
For example, for the instance in Fig. 8, $a_{2}$ and $a_{3}$ are unavoidable agents of $a_{1}$, as $a_{1}$’s dependence path must pass through $T[a_{2}]$ and $T[a_{3}]$.
Definition 10.
Unavoidance graph $\mathcal{G}_{u}$ of a cluster: an undirected representation of whether one agent is unavoidable to another agent within the cluster. Assume there are two agent $a_{i}$ and $a_{j}$, $a_{i}$ is unavoidable to $a_{j}$ if $a_{j}$’s dependence path must pass $a_{i}$’s start or target, meanwhile all dependence paths cannot pass a agent’s start or target if the agent is not in current cluster. It is important to note that for a given cluster $r$, its unavoidance graph is unique.
An example of $\mathcal{G}_{u}$ can be found in Fig. 10.
@@
Intuitively, these maximum unavoidable agents represent the largest undividable subset within the current cluster, are referred to as “major set" during the cluster bipartition process. Agents within the cluster, excluding those belonging to the maximum unavoidable agents, are referred to as “remaining set" during the cluster bipartition process. An example of the maximum unavoidable agents of a cluster can be found in Fig. 10.
It is noteworthy that during the bipartition process, agents in the remaining set will be moved to the major set, but agents in the remaining set will not move into the major set until both the major set and the remaining set are independent.
@@
1:$r,\mathcal{G}_{c}$
$\mathcal{L}$
3:///// / determine initial dependence paths
4:$P=\emptyset$; ///// / $P[a]$ means agent a’s denpendence path
5:for agent $a$ in $r$ do
6:     $p_{i}=search\_path\_SAT(a_{i},\mathcal{G}_{c},r,\emptyset)$;
7:     add $p_{i}$ to $P$;
8:end for
9:get $\mathcal{G}_{s}$ from $P$;
10:determine all loop in $\mathcal{G}_{s}$ via strong component detect;
11:///// / sort levels;
12:construct graph $\mathcal{G}_{l}$ via $\mathcal{G}_{s}$;
13:$root\_levels$ = level in $\mathcal{G}_{l}$ that not later than other level;
14:set order of all level to 0;
15:$levels$ = $root\_levels$;
16:while levels is not empty do
17:     $next\_levels$ = $\emptyset$;
18:     for level in levels do
19:         for another_level that later than level in $\mathcal{G}_{l}$ do
20:              if $another\_level$’s order $\leq$ level’s order  then
21:                  $another\_level$’s order = level’s order + 1;
22:                  add $another\_level$ to $next\_levels$;
23:              end if
24:         end for
25:     end for
26:     $levels$ = $next\_levels$;
27:end while
28:return all levels in the order of last visit to them;
@@
1, identify the maximum unavoidable agents of the cluster, referred to as the major set (Line 1 $\sim$ 12 in Algorithm 2);
2, examine each agent in the remaining set to determine if its dependence path must pass through agents of the major set. Move agents meeting this criterion to the major set, until the remain set is independent or empty(Line 16 $\sim$ 27 in Algorithm 2);
3, check any newly added agents to the major set to ascertain if their dependence paths must pass through agents in the remaining set. Transfer such agents from agents the remaining set to the major set, until the major set is independent(Line 29 $\sim$ 40 in Algorithm 2);
4, verify whether the major set and the remaining set both meet the legality requirements of the cluster (for every agent in a cluster, its dependence path does not pass through the start or target states of agents in other clusters). If they do (or if the remaining set is empty), exit and return the major set and the remaining set as the result of the bipartition. (Line 42 $\sim$ 44 in Algorithm 2) Otherwise, proceed to step 2;
@@
The pseudocode for bipartitioning a cluster is outlined in Algorithm 2. An example of the bipartition of a cluster is shown in Fig. 10. The verification of the legality of these clusters is shown in Fig. 11.
@@
1:$\{P_{1},...,P_{m}\}$ ///// / assume there are $m$ subproblems and their solution is $P_{i}$
2:$\{P_{1},...,P_{m}\}$
3:set last occupied time of all cells to 0;
4:for each solution $P$ in $\{P_{1},...,P_{m}\}$ do
5:     while True do
6:         $t$=1; ///// / current time index
7:         ///// / check whether insert wait action to $P$
8:         $need\_wait$ = False;
9:         $all\_finished$ = True;
10:         for path $p$ in $P$ do
11:              if t $>$ size of $p-1$ then
12:                  ///// / don’t check path that have reach target
13:                  continue;
14:              else
15:                  $all\_finished$ = False;
16:              end if
17:              ///// / $p[t]$ means path $p$’s location at time $t$
18:              if last occupied time of $p[t]\geq t$ then
19:                  $need\_wait$ = True;
20:                  break;
21:              end if
22:         end for
23:         if all_finished then
24:              break;
25:         end if
26:         ///// / insert wait action to all path in $P$
27:         if need_wait then
28:              for path $p$ in $P$ do
29:                  if t $>$ size of $p-1$ then
30:                       continue;
31:                  end if
32:                  repeat $p[t]$ in $p$ at time index $t$;
33:              end for
34:         else
35:              ///// / update last occupied time of cells
36:              for path $p$ in $P$ do
37:                  if t $>$ size of $p-1$ then
38:                       continue;
39:                  end if
40:                  set last occupied time of $p[t]$ to $t$;
41:              end for
42:         end if
43:         $t$ = $t$ + 1;
44:     end while
45:end for
46:return $\{P_{1},...,P_{m}\}$;